OpenStudy (anonymous):

Anyone a genius in physics? Which of the following systems reflects a decrease in entropy? burning match freezing water evaporating alcohol breaking glass

OpenStudy (anonymous):

i got evaporating alcohol

OpenStudy (anonymous):

@bibby do you know physics?

OpenStudy (anonymous):

@kidrah69 yo poojah! you know physics dude?

OpenStudy (anonymous):

i just need someone to clarify that i got this right.

OpenStudy (anonymous):

freezing water

OpenStudy (anonymous):

@icecreamman

OpenStudy (anonymous):

yeah it is the water freeze, when water freezes it is changed from liquid state (more entropy) to solid state ( less entropy) thus we have a decrease in entropy

OpenStudy (shamim):

For measuring entropy, the formula is dS=dQ/T

OpenStudy (shamim):

Here dS=change of entropy dQ=change of heat taken by the the system T=temperature in kelvin scale

OpenStudy (shamim):

Just c dQ will b negative

OpenStudy (shamim):

Because dQ=Q2-Q1=something negative

OpenStudy (shamim):

Q2=heat of freeze Q1=heat of hot water

OpenStudy (shamim):

So dS will b negative

OpenStudy (shamim):

So change of entropy is negative??????!!!!!

OpenStudy (shamim):

Response plz!!!!!!!

OpenStudy (anonymous):

Sorry i was out but freezing water? Why isnt it evap alcohol?

OpenStudy (shamim):

Did u c my equation for measuring change of entropy

OpenStudy (shamim):

For evaporation of alcohol dQ is positive So dS is also positive So entropy increasing Right????!!!

OpenStudy (shamim):

T can not b negative

OpenStudy (shamim):

T= tamperature in kelvin scale

OpenStudy (anonymous):

ooooooohhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhh!!!!!!!!!!!! i get it!!! thank you so much!!!!!!!!!!!!

OpenStudy (shamim):

U r most welcomeeeee!!!!!!!!!!!

OpenStudy (anonymous):

@shamim can you help with another question that i have?